a cars velocity decreases uniformly 20 m/s to 6m/s while covering 70m distance in a straight line. Find its acceleration and time taken to the travel distance



1
Expert's answer
2022-06-09T09:59:14-0400

Given:

"v_i=20\:\rm m/s"

"v_f=6\:\rm m/s"

"d=70\:\rm m"



"d=\frac{v_f^2-v_i^2}{2a}"

"a=\frac{v_f^2-v_i^2}{2d}=\frac{6^2-20^2}{2*70}=-2.6\:\rm m/s^2"